OpenStudy (anonymous):

solve the following system of equations. -1/2x+1/5y=3 2x+1/3y=-2/3

OpenStudy (anonymous):

can u rewrite the first equation ?it is not clear ?!

OpenStudy (anonymous):

\[\frac{ ?1 }{ ? 2}x+\frac{ ? 1}{ ?5 }y=3\]

OpenStudy (anonymous):

minus the question marks ofc

OpenStudy (anonymous):

\[2x+\frac{ 1 }{ 3 }y=-\frac{ 2 }{ 3}\]